# Polymer Collapse & Liquid-Liquid Phase-Separation are Coupled in a Generalized Prewetting Transition

### v1

@@ Results
Strongly interacting bulk drives a first-order transition — In the limit where the bulk system is instead near to phase coexistence, when $t_{\text{b}}<0$ the bulk free energy density (equation 5) is not well described by a single quadratic minimum. Still, the order parameters $R_{\text{g}}$ and $\phi$ take values which minimize the combined free energy $F_{\text{sys}}$ (equation 7). While this minimization cannot be done analytically, minima are easy to find numerically. Contour plots of $F_{\text{sys}}$ are shown in Figure 3D,E (with $\rho$ rather than $R_{\text{g}}$ on the x axis), where there are two distinct minima. As parameters change, there is an abrupt transition as one minimum becomes lower in free energy than the other. This leads to discontinuities in the polymer density (blue curve in Figure  3B) and in the bulk order parameter (blue curve in Figure  3C).
Figure 4: Bulk proteins change the configuration of a multi-component polymer — (A) (upper) Illustration of a multi-component polymer: the polymer has discrete, contiguous segments of monomers that interact differently with the bulk molecules. Here yellow interacts strongly while green is non-interacting. (lower) Schematic showing the selective collapse of yellow segments into a Prewet phase. (B) Occupancy of yellow and green monomers for the bulk binding molecules, as a function of $\mu_{\text{b}}$. Collapse of the yellow segments coincides a jump in the occupancy of yellow segments, but not green. Letters in the plot correspond to the figures below. (C-E) Simulation configurations, monomer contact probabilities, and occupancy profiles of different multi-component polymer phases. (C) Extended phase: the polymer is Extended and bulk occupancy is low and near-uniform. (D) In non-interacting bulk systems, the occupancy of the yellow regions is high (bottom), but the polymer remains in an Extended configuration (middle). (E) Global collapse: (middle) long range, inter-segment contacts are formed via interactions by the bulk phase, red off-diagonal regions in the contact probabilities plot. (lower) Both red and blue bulk molecules are enriched in the yellow segments, but not in the green.
@@
To gain additional insight into the nature of this abrupt transition, we can also approximate the free energy as quadratic around both dense and dilute phase minima, treating each with the same approximation as above (orange and purple lines in Figure 3D,E left). This predicts that local minima will lie near the corresponding orange and purple lines in Figure  3D,E, which meet the y-axis at the locations of the local minima of the bulk free energy density. In this approximation, the system takes the position of either the purple or orange dot, and a first order transition occurs when they have the same value of $F_{\text{sys}}$. The free energy about the bulk dense phase sees an additional contribution proportional to $R^{3}\mu_{\text{b}}\phi_{\infty}$. This penalizes Extended configurations and drives the polymer to collapse in an abrupt prewetting transition that accompanies moving to the local minimum associated with the bulk dense phase. Our simulations and theory show that $\rho\rightarrow 0$ as $\mu_{\text{b}}\rightarrow 0$. This suggests that, in the thermodynamic limit, the Wet phase corresponds to an Extended polymer localized to the bulk dense phase as long as $h>0$.
@@
In many biological scenarios bulk proteins interact strongly with certain regions of the chromosome, and weakly or not at all with others. We modified our simulations by partitioning the long polymer into discrete segments that either interact with the bulk (yellow in Figure 4A) or do not (green in Figure 4A). We use this to understand how polymer sequence interacts with the behavior we describe above.
Since the bulk interacts differently with the two monomer types, there are situations where strongly interacting segments collapse but non-interacting segments remain Extended. At low $\mu_{\text{b}}$ the polymer is Extended (Figure  4C) while at high $\mu_{\text{b}}$ all interacting segments form a globally Collapsed phase (Figure 4E). These phase differ in their occupation of bulk molecules and their polymer-polymer contacts. The Extended phase sees low occupancy of directly binding bulk molecules (red) at the yellow segments (4C, bottom) while both direct and indirect (blue) bulk molecules are enriched in the globally Collapsed phase  4E, bottom). The occupancy of yellow monomers increases discontinuously with $\mu_{\text{b}}$ while the green monomers remain mostly unoccupied (Figure  4B). When bulk molecules do not interact with each other the occupancy of yellow regions increases smoothly with $\mu_{\text{bulk}}$ (4B, diamonds), and there is no accompanying occupancy in indirect binders (4D, lower).
Polymer-polymer contacts in the Extended phase are homogeneous, while the globally Collapsed phase sees long-range contacts between monomers of the same type, even those separated by an intervening sequence (4E,middle). In this globally Collapsed phase the reduction in surface tension associated with merging domains must overcome the entropic penalty of looping non-interacting stretches as occurs in our simulations. We also see transient but long-lived locally Collapsed states, and there may be regimes where multiple identical domains stably coexist without merging.
@@ Discussion
Transcriptional activation through a prewetting transition — Transcription factors interact directly with enhancers, and many feature ‘activation domains’ that phase-separate with transcriptional machinery such as the co-activator Mediator [23, 7, 24], playing roles roughly analogous to the red and blue polymers in our simulations. These Prewet phases could naturally integrate information from transcription factor concentrations into transcriptional regulation by controlling where along the chromosome Prewet domains occur (Figure  4B). Transcription factor concentration could also modify how enhancers, promoters, and distant genomic regions interact in three dimensions, as in Figure 4E,F. Indeed, many Eukaryotic promoters become spatially proximal to regulatory elements during transcription initiation [25, 26, 2, 3].
Transcriptional repression with Prewet phases — Many proteins that interact with heterochromatin and other transcriptionally repressed regions have a propensity to phase-separate in the presence or absence of DNA [8, 9, 27, 28, 29]. Heterochromatin is typically more dense than transcriptionally active chromatin, and is inaccessible to transcriptional machinery [30]. It is possible that heterochromatin is best thought of as a separate generalized Prewet phase, which excludes transcriptional machinery, and is rich in DNA and certain proteins. The complex interactions between the sequence of regulatory elements and cell-type specific transcription factors could regulate transcription by determining which regions are prewet by transcriptionally inaccessible phases [28, 29].
Prewet phases permit specificity in transcriptional regulation — Eukaryotes face enormous genetic ‘crosstalk’  [31] in that many transcription factors bind promiscuously to many potential genomic targets, so that each individual binding site on the genome carries little specificity. The prewetting phase transition naturally integrates over a long stretch of chromosome, yielding an effective cooperativity in when and where domains appear. While the implications of regulation in this manner is not our focus here, we do expect that long stretches of chromosome with many weak transcription factor binding sites could undergo collapse transitions in a manner that is highly sensitive to small changes in transcription factor concentrations.
@@
The composition and genomic localization of Prewet domains can encode and read out cellular identity — Cellular identity is thought to be encoded by core regulatory complexes (CoRCs) composed of a small number of transcription factors which regulate themselves and downstream cell-type specific genes [32, 33]. Evidence for these complexes comes in part from probes of DNA-Protein interactions. Chromatin Immuno-precipitation (ChIP-seq) studies probe the spatial proximity of specific proteins to specific regions of the genome, and often find long stretches enriched in cell-type specific groups of a few transcription factors. Proteins that map to the same regions are often interpreted as arising from large, macromolecular complexes. But a generalized Prewet phase localized to certain genomic regions has a similar experimental signature, as Figure  4E (occupancy) illustrates. Molecules that do not bind directly to the polymer (blue) localize to the same region of the polymer as molecules that have direct interactions (red) so long as both of these molecules phase-separate with each other. Many transcription factors that inspired this model for CoRCs phase-separate at high concentrations (OCT4, NANOG  [23]), and are found as large, localized assemblies in some cell types via microscopy. Prewet phases which depend on cellular identity may also explain cell type specific differences in the three-dimensional structure of chromatin (Figure  4E, polymer contacts).
@@
Phase diagrams for generalized Prewet phases — Past work has investigated how binary bulk mixtures can lead to polymer collapse even when either component on its own does not, a phenomena known as polymer-co-nonsolvency [34]. As in that work, far above the critical point in our more complicated bulk mixture, the transition from an Extended to a Collapsed state is continuous, though steep. Near the two phase coexistence region of a complicated bulk this transition becomes abrupt, as we see in simulations and mean-field theory. This implies a tri-critical point which we have not thoroughly explored, near but outside of bulk coexistence, where the line of first-order prewetting transitions meets the line of second order transitions. By analogy to prewetting to surfaces, we also anticipate that the line of prewetting transitions meets bulk coexistence at another tricritical point, where Dry, Wet, and Prewet phases coexist. At these tricritical points the polymer likely sees non-trivial scaling behavior. We also expect interesting finite-size effects in the wetting regime when the length of the polymer is such that its Collapsed size is of order the size of the bulk domain. And finally, we anticipate coupling many-component bulk systems to a multi-component polymer will yield very rich physics [35].
@@
Our model differs from past models of phase separation and chromosomes — Recent work has investigated the role of polymer ‘scaffolds’ on the phase-separation of $3D$ proteins [36, 37, 10], and how $3D$ proteins mediate communication between distal genomic regions [38, 39]. This work has emphasized that polymers can widen regimes in which a bulk fluid undergoes a phase-transition. Our results are also related to a body of work investigating how demixing transitions interact with polymer meshes [40, 41], though most of that literature has focused on cytoskeletal meshes, which interact elastically, but do not contribute a transition of their own. Depending on the sign and strength of the interaction between liquid droplets and mesh components, demixing transitions can be enhanced or inhibited, and are generically accompanied by distortions of the polymer mesh. Here we innovate by jointly considering the polymer collapse transition and bulk condensation. Our work is also related to models of the ParABS system for chromosome segregation in bacteria, which explicitly describe the interactions of a dynamic polymer with interacting proteins, but do not focus on the phase behavior of their bulk system [42].
@@
Distinguishing prewetting from wetting — Although we focus on a region of prewetting phase transitions, our model features other transitions that resemble wetting, where a bulk phase prefers to localize to a surface [16]. Wet phases are stable even without a long polymer; while polymers can partition into them, the presence of a Wet phase does not imply presence of a specific polymer. As such, Wet phases are less suitable for regulation by controlling DNA’s interaction with specific components or its three dimensional structure. Overexpression of transcriptional activators quite generally inhibits transcription through a phenomenon dubbed ‘squelching’ [43]. It is possible that squelching marks a transition into bulk coexistence where macroscopic droplets compete with the DNA for transcriptional machinery.
@@
Other types of generalized prewetting may be common — This work and a similar model we proposed to explain membrane-localized signaling clusters  [15] contain two defining features: (1) a bulk fluid that de-mixes in $3D$, and (2) a lower-dimensional surface with a phase transition that is different from that of the bulk. In both of these cases the phase-behavior of the surface gives the bulk qualities of a transition that it didn’t have before – a $2D$ critical point in the membranes case, and qualities of a polymer collapse transition here. Eukaryotic cells have many surfaces with phase-transitions that do not occur in bulk, and biology often needs to localize processes to specific locations. We speculate that biology uses a variety of these generalized-prewetting transitions to accomplish specific localization, and to imbue a bulk fluid with properties such as a diverging susceptibility. Work by ourselves and others [44, 45, 41, 10] will continue to uncover the rich physics allowed by surface phase-transitions.

### v2

@@ Abstract
Strongly interacting bulk drives a first-order transition — In the limit where the bulk system is instead near to phase coexistence, when $t_{\text{b}}<0$ the bulk free energy density (equation 5) is not well described by a single quadratic minimum. Still, the order parameters $R_{\text{g}}$ and $\phi$ take values which minimize the combined free energy $F_{\text{sys}}$ (equation 7). Contour plots of $F_{\text{sys}}$ are shown in Fig.3D,E (with $\rho$ rather than $R_{\text{g}}$ on the x axis), where there are two distinct minima. When the long polymer interacts more strongly with itself, (lower $v_{\text{p}}$) there is an abrupt transition as one minimum becomes lower in free energy than the other. This leads to discontinuities in the polymer density (blue curve in Fig.3B) and in the bulk order parameter (blue curve in Fig.3C).
Figure 4: Lattice monte-carlo simulations confirm mean-field theory predictions: — (A) Density difference between molecules on the polymer and in the bulk as function of $\mu_{\text{bulk}}$, for three values of $J_{\text{bulk}}$. For $J_{\text{bulk}}=1.0,1.5$ there is a discontinuous change in the bulk density difference as $\mu_{\text{bulk}}$ is increased towards coexistence, and an analogous first order transition as the bulk enters its dense phase. For $J_{\text{bulk}}=0.5$ there is a steep but continuous increase in the bulk density difference through the transition, and a corresponding transition as the polymer de-condenses; Cf. 3C. (B) The polymer collapse transition shifts and becomes discontinuous in an interacting solvent. Cf 3B. (C) Simulation configurations corresponding to blue and purple circles in (A). For all simulation parameters, see Supplementary Material
@@
To gain additional insight into the nature of this abrupt transition, we can also approximate the free energy as quadratic around both dense and dilute phase minima, treating each with the same approximation as above (orange and purple lines in Fig 3D,E left). This predicts that local minima will lie near the corresponding orange and purple lines in Fig  3D,E. In this approximation, the system takes the position of either the purple or orange dot, and a first order transition occurs when they have the same value of $F_{\text{sys}}$. Our simulations and theory show that $\rho\rightarrow 0$ as $\mu_{\text{b}}\rightarrow 0$, corresponding to an Extended polymer wet by the bulk dense phase.
Polymer collapse can only be accessed from the extended phase through a phase transition, either continuous or abrupt. Our simulations are in the abrupt regime, but the presence of a continuous transition above $T_{\text{c}}$, suggests that a generic phase diagram has the features plotted in figure 2E; a line of abrupt transitions in the dilute bulk phase but just outside of bulk coexistence, a line of continuous transitions at hight $T_{\text{b}}$ and near the bulk dense phase, and a tricritical point where they meet. Near the bulk critical point but in the dense phase, a narrow region of prewetting can be accessed through a continuous transition by decreasing bulk density, and thereby increasing effective interactions. This is related to the phenomenon of polymer co-nonsolvency [25], in which a mixture of a good and bad solvent can induce polymer collapse more effectively than the bad solvent in isolation.
Monte-Carlo simulations support theoretical predictions — Some predictions from our mean-field theory are testable in our Monte-Carlo simulations : (i) Is there a regime of continuous transitions at weak $J_{\text{bulk}}$, and a regime of discontinuous transitions at large $J_{\text{bulk}}$? (ii) Do interactions from the bulk solvent shift the collapse transition of the long polymer. We tested these predictions in a minimal Lattice Monte-Carlo Simulation (see Methods). Figure 4A demonstrates a regime where polymer and bulk density increase abruptly as $\mu_{\text{b}}$ increases for $J_{\text{bulk}}>>0$ (blue line), and a regime where the densities continuously change for weaker $J_{\text{bulk}}$. (purple line). Likewise Figure 4B supports the ’shifting’ of the polymer collapse transition with solvent-mediated interactions. This shift can be continuous for weak coupling strengths (purple line) or discontinuous for stronger couplings (blue line).
The discontinuities we predict will be blunted in finite-size systems. The sharpness of the prewetting-collapse transition is set by the contour length of the long polymer $N_{\text{p}}$. While biological polymers are larger than most synthetic polymers, a direct mapping between our $N_{\text{p}}$ and monomer units in biological systems is unclear. Finite size bulk systems can also blunt these discontinuities but only through material depletion, an effect which we do not consider.
@@
Figure 5: Bulk proteins change the configuration of a multi-component polymer — (A) (upper) Illustration of a multi-component polymer: the polymer has discrete, contiguous segments of monomers that interact differently with the bulk molecules. Here yellow interacts strongly while green is non-interacting. (lower) Schematic showing the selective collapse of yellow segments into a Prewet phase. (B) Occupancy of yellow and green monomers for the bulk binding molecules, as a function of $\mu_{\text{b}}$. Collapse of the yellow segments coincides a jump in the occupancy of yellow segments, but not green. Letters in the plot correspond to the figures below. (C-E) Simulation configurations, monomer contact probabilities, and occupancy profiles of different multi-component polymer phases. For simulation parameters, see SI (C) Extended phase: the polymer is Extended and bulk occupancy is low and near-uniform. (D) In non-interacting bulk systems, the occupancy of the yellow regions is high (bottom), but the polymer remains in an Extended configuration (middle). (E) Global collapse: (middle) long range, inter-segment contacts are formed via interactions by the bulk phase, red off-diagonal regions in the contact probabilities plot. (lower) Both red and blue bulk molecules are enriched in the yellow segments, but not in the green.
In many biological scenarios bulk proteins interact strongly with certain regions of the chromosome, and weakly or not at all with others. We modified our simulations by partitioning the long polymer into discrete segments that either interact with the bulk (yellow in Fig 5A) or do not (green in Fig 5A).
Since the bulk interacts differently with the two monomer types, there are situations where strongly interacting segments collapse but non-interacting segments remain Extended. At low $\mu_{\text{b}}$ the polymer is Extended (Fig  5C) while at high $\mu_{\text{b}}$ all interacting segments form a globally Collapsed phase (Fig 5E). These phases differ in their occupation of bulk molecules and their polymer-polymer contacts. The Extended phase sees low occupancy of directly binding bulk molecules (red) at the yellow segments (Fig 5C, bottom) while both direct and indirect (blue) bulk molecules are enriched in the globally Collapsed phase (Fig  5E, bottom). The occupancy of yellow monomers increases discontinuously with $\mu_{\text{b}}$ while the green monomers remain mostly unoccupied (Fig  5B). When bulk molecules do not interact with each other, $J_{\text{bulk}}=0$, the occupancy of yellow regions increases smoothly with $\mu_{\text{b}}$ (Fig 5B, diamonds), and there is no accompanying occupancy in indirect binders (Fig 5D, lower). Polymer-polymer contacts in the Extended phase are homogeneous, while the globally Collapsed phase sees long-range contacts between monomers of the same type, even those separated by an intervening sequence (Fig 5E,middle).
@@ Discussion
Transcriptional activation through a prewetting transition — Transcription factors interact directly with enhancers, and many feature ‘activation domains’ that phase-separate with transcriptional machinery such as the co-activator Mediator [26, 7, 27], playing roles roughly analogous to the red and blue polymers in our simulations. These Prewet phases could naturally integrate information from transcription factor concentrations into transcriptional regulation by controlling where along the chromosome Prewet domains occur (Fig  5B). Transcription factor concentration could also modify how enhancers, promoters, and distant genomic regions interact in three dimensions, as in Fig 5E,F. The prewetting phase transition naturally integrates over a long stretch of chromosome, yielding an effective cooperativity in when and where domains appear, as may be relevant for transcription initiation [28, 29, 2, 3]. It is possible that heterochromatin is best thought of as a separate generalized Prewet phase, which excludes transcriptional machinery, and is rich in DNA and certain proteins. [8, 10, 30, 31, 32]
@@
The composition and genomic localization of Prewet domains can encode cellular identity — Cellular identity is thought to be encoded by core regulatory complexes (CoRCs) composed of a small number of transcription factors which regulate downstream cell-type specific genes [33, 34]. Evidence for these complexes comes, in part, from Chromatin Immuno-precipitation (ChIP-seq) studies [35]. Long stretches enriched in cell-type specific groups of a few transcription factors are often interpreted as arising from large, macromolecular complexes. But a generalized Prewet phase localized to certain genomic regions has a similar experimental signature, as Fig.5E (occupancy) illustrates. Molecules that do not bind directly to the polymer (blue) localize to the same region as molecules that interact directly (red) so long as both of these molecules phase-separate with each other. Many transcription factors that inspired this model for CoRCs phase-separate at high concentrations (OCT4, NANOG  [26]), and are found as large, localized assemblies in some cell types. Prewet phases which depend on cellular identity may also explain cell type specific differences in the three-dimensional structure of chromatin (Fig.5E, polymer contacts).
@@
Phase diagrams for generalized Prewet phases — Past work has investigated how binary bulk mixtures can lead to polymer collapse even when either component on its own does not, a phenomena known as polymer-co-nonsolvency [25]. As in that work, far above the critical point in our more complicated bulk mixture, the transition from an Extended to a Collapsed state is continuous, though steep. Near the two phase coexistence region this transition becomes abrupt, as we see in simulations and mean-field theory  3A, compare blue and purple lines. This implies a tri-critical point which we have not thoroughly explored, near but outside of bulk coexistence, where the line of first-order prewetting transitions meets the line of second-order transitions. We also anticipate that the line of prewetting transitions meets bulk coexistence at another tricritical point, where Dry, Wet, and Prewet phases coexist. At these tricritical points the polymer likely sees non-trivial scaling behavior. We also expect interesting finite-size effects in the wetting regime when the length of the polymer is such that its Collapsed size is of order the size of the bulk domain. And finally, we anticipate coupling many-component bulk systems to a multi-component polymer will yield very rich physics [36].
@@
Our model differs from past models of phase separation and chromosomes — Recent work has investigated the role of polymer ‘scaffolds’ on the phase-separation of $3D$ proteins [37, 38, 11, 39, 40, 15], and how $3D$ proteins mediate communication between distal genomic regions [41, 42, 9]. Others have observed that long polymers can widen regimes in which bulk proteins phase separate; here we expand on these observations by studying the thermodynamics of these transitions Our results are also related to a body of work investigating how demixing transitions interact with polymer meshes [43, 44], though most of that literature has focused on cytoskeletal meshes, which interact elastically, but do not contribute a transition of their own. There are regimes where demixing transitions can be enhanced or inhibited, and are generically accompanied by distortions of the polymer mesh. Here we innovate by jointly considering the polymer collapse transition and bulk condensation. Our work is also related to models of the ParABS system for chromosome segregation in bacteria, which explicitly describe the interactions of a dynamic polymer with interacting proteins, but do not focus on the phase behavior of their bulk system [45].
@@
Distinguishing prewetting from wetting — Although we focus on a region of prewetting phase transitions, our model features other transitions that resemble wetting, where a bulk phase prefers to localize to a surface [18]. Wet phases are stable even without a long polymer; while polymers can partition into them, the presence of a Wet phase does not imply presence of a specific polymer. As such, Wet phases are less suitable for regulation by controlling DNA’s interaction with specific components or its three dimensional structure. Overexpression of transcriptional activators quite generally inhibits transcription through a phenomenon dubbed ‘squelching’ [46]. It is possible that squelching marks a transition into bulk coexistence where macroscopic droplets compete with the DNA for transcriptional machinery. A complementary view is that chromatin-localized condensates serve as heterogeneous nucleation seeds for specific bulk phases [15]. We interpret these ’seeds’ as bona-fide prewet phases that may interplay with wet phases in genetic regulation.
@@
Eukaryotic cells have many surfaces with phase-transitions that do not occur in bulk, and biology often needs to localize processes to specific locations. We speculate that biology uses a variety of these generalized-prewetting transitions to accomplish specific localization, and to imbue a bulk fluid with properties such as a diverging susceptibility. Work by ourselves and others  [47, 48, 44, 11] will continue to uncover the rich physics allowed by surface phase-transitions.
Author Contributions — M.N.R. and B.B.M. conceived the project. M.N.R carried out the research. M.N.R and B.B.B. analyzed the data and wrote the paper.
@@
Declaration of Interests — The authors declare no competing interests.